### v3

# Towards Understanding Link Predictor Generalizability Under Distribution Shifts

Abstract.
State-of-the-art link prediction (LP) models demonstrate impressive benchmark results. However, popular benchmark datasets often assume that training, validation, and testing samples are representative of the overall dataset distribution. In real-world situations, this assumption is often incorrect; uncontrolled factors lead new dataset samples to come from a different distribution than training samples. Additionally, the majority of recent work with graph dataset shift focuses on node- and graph-level tasks, largely ignoring link-level tasks. To bridge this gap, we introduce a novel splitting strategy, known as LPShift, which utilizes structural properties to induce a controlled distribution shift. We verify LPShift’s effect through empirical evaluation of SOTA LP models on 16 LPShift variants of original dataset splits, with results indicating drastic changes to model performance. Additional experiments demonstrate graph structure has a strong influence on the success of current generalization methods. The source code is available here https://github.com/revolins/LPShift.
Link Prediction, Synthetic Graphs, Distribution Shift, Graph Neural Networks
journalyear: 2025copyright: ccconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737434isbn: 979-8-4007-1454-2/2025/08ccs: Mathematics of computing Graph algorithmsccs: Computing methodologies Supervised learning
1. Introduction
Link Prediction (LP) is concerned with predicting unseen links (i.e., edges) between two nodes in a graph (Liben-Nowell and Kleinberg, 2003). The task has a wide variety of applications including: recommender systems, (Fan et al., 2019), knowledge graph completion (Lin et al., 2015), protein-interaction (Kovács et al., 2019), and drug discovery (Abbas et al., 2021). Traditionally, LP was performed using heuristics that model the pairwise interaction between two nodes (Newman, 2001; Zhou et al., 2009; Adamic and Adar, 2003). The success of Graph Neural Networks (GNNs) (Kipf and Welling, 2017) has prompted their usage in LP (Kipf and Welling, 2016; Zhang and Chen, 2018). However, GNNs are unable to fully-capture representations for node pairs (Zhang et al., 2021; Srinivasan and Ribeiro, 2019). To combat this problem, recent methods (i.e., GNN4LP) empower GNNs with additional information to capture pairwise interactions between nodes (Shomer et al., 2024; Chamberlain et al., 2022; Zhang and Chen, 2018; Wang et al., 2023a) and demonstrate tremendous ability to model LP on real-world datasets (Hu et al., 2020).
While recent methods have shown promise, current benchmarks (Hu et al., 2020) typically assume that the training and evaluation data is drawn from the same structural distribution. This assumption collapses in real-world scenarios, where the structural feature (i.e., covariate) distribution may shift from training to evaluation. Therefore, it’s often necessary for models to generalize to samples whose newly-introduced feature distribution differs from the training dataset (Wiles et al., 2021; Yao et al., 2022a, b; Bevilacqua et al., 2021).
Furthermore, while numerous methods work to account for distribution shifts within graph machine learning (Li et al., 2022b), there remains little work doing so for LP. Specifically, we observe that (1) No LP Benchmark Datasets: Current graph benchmark datasets designed with a quantifiable distribution shift are focused solely on the node and graph tasks (Zhou et al., 2022a; Ding et al., 2021), with no datasets available for LP. (2) Absence of Foundational Work: There is limited existing work for distribution shifts relevant to LP (Zhang et al., 2022; Dong et al., 2022). Current methods are primarily focused on detecting and alleviating anomalies within node- and graph-level tasks (Jin et al., 2022; Bevilacqua et al., 2021; Gao et al., 2023; Wu et al., 2024; Guo et al., 2023; Wu et al., 2023; Sui et al., 2024; Li et al., 2022b). Additionally, few methods are designed for aiding LP generalization in any setting (Dong et al., 2022; Zhao et al., 2022; Zhang et al., 2022; Zhou et al., 2022a). Also, other LP generalization methods which are theorized to improve performance in shifted scenarios remain crucially untested (Singh et al., 2021; Wang et al., 2023b).
To tackle these problems, this work proposes the following contributions: (1) Creating Datasets with Meaningful Distribution Shifts. LP requires pairwise structural considerations (Liben-Nowell and Kleinberg, 2003; Mao et al., 2024a). Additionally, when considering realistic settings (Li et al., 2024) or distribution shift (Zhu et al., 2024), GNN4LP models are sensitive to structual changes within datasets relative to graph (Wei et al., 2021; Yuan et al., 2021) and node classification (Shi et al., 2023; Zhao et al., 2023) models. To quantify this sensitivity and better understand distribution shifts, we use key structural LP heuristics to split the links into train/validation/test splits via LPShift. By applying LPShift to generate dataset splits, we induce shifts in the underlying distribution of the links, thereby affecting link formation (Mao et al., 2024a) and downstream performance. This elevates preexisting LP benchmarks by allowing users to control the extent with which their dataset undergoes distribution shift while retaining relevance to LP. Much like how prior benchmarks induce substructure (Gui et al., 2022; Koh et al., 2021) and drug scaffold (Ji et al., 2022) shift to provide relevant challenges within graph and node classification. Theoretical justification is provided in Section 4 and empirical results are provided in Section 5. (2) Benchmarking Current LP Methods. GNN4LP models demonstrate a strong sensitivity to LPShift when generalizing to the synthetic dataset splits. Despite the existence of LP generalization methods, such as FakeEdge (Dong et al., 2022) and Edge Proposal Sets (Singh et al., 2021), there remains little work benchmarking link-prediction models under distribution shifts (Ding et al., 2021; Dong et al., 2022; Zhu et al., 2024). This lack of benchmarking contributes to a gap in understanding, impeding the capabilities of LP models to generalize in real-world scenarios. This work quantifies the performance of SOTA LP models under 16 unique LPShift scenarios and provides analysis as a foundation for improving LP model generalization. We further quantify the effects of LP and graph-specific generalization methods, finding that they also struggle to generalize with differing structural shifts.
The remainder of this paper is structured as follows. In Section 2, we provide background on the heuristics, models, and generalization methods used in LP. In Sections 3 and 4, we detail how structural heuristics, theoretical perspectives, and HeaRT (Li et al., 2024) relate to our proposed splitting strategy. After which, we formally introduce LPShift. Lastly, in Section 5, we benchmark a selection of LP models and generalization methods on LPShift, followed by analysis to understand the effects of this new strategy.
2. Related Work
LP Heuristics: Classically, neighborhood heuristics, which measure characteristics between source and target edges, functioned as the primary means of predicting links. These heuristics show limited effectiveness with a relatively-high variability in results, largely due to the complicated irregularity within graph datasets; which only grows worse with larger datasets (Liben-Nowell and Kleinberg, 2003). Regardless of this, state-of-the-art GNN4LP models have integrated these neighborhood heuristics into neural architectures to elevate link prediction capabilities (Wang et al., 2023a; Chamberlain et al., 2022).
For a given heuristic function, $u$ and $v$ represent the source and target nodes in a potential link, $(u,v)$. $\mathcal{N}(v)$ is the set of all edges, or neighbors, connected to node $v$. $f(v_{i,i+1},u)$ is a function that considers all paths of length $i$ that start at $v$ and connect to $u$. The three tested heuristics are as follows:
Common Neighbors (Newman, 2001): The number of neighbors shared by two nodes $u$ and $v$,
(1) $\text{CN}(u,v)=|\mathcal{N}(u)\cap\mathcal{N}(v)|.$
Preferential Attachment (Liben-Nowell and Kleinberg, 2003): The product of the number of neighbors (i.e., the degree) for nodes $u$ and $v$,
(2) $\text{PA}(u,v)=|\mathcal{N}(u)|\times|\mathcal{N}(v)|.$
Shortest Path Length (Liben-Nowell and Kleinberg, 2003): The path between $u$ and $v$ which considers the smallest possible number of nodes, denoted as length $n$,
(3) $\text{SP}(u,v)=\operatorname*{arg\,min}_{\Sigma}(\Sigma^{n-1}_{i=1}f(v_{i,i+1},u)).$
GNNs for Link Prediction (GNN4LP): LP’s current SOTA methods rely on GNNs for a given model’s backbone. The most common choice is the Graph Convolutional Network (GCN) (Kipf and Welling, 2017), integrating a simplified convolution operator to consider a node’s multi-hop neighborhood. The final score (i.e., probability) of a link existing considers the representation between both nodes of interest. However, (Zhang et al., 2021) show that such methods aren’t suitably expressive for LP, as they ignore vital pairwise information that exists between both nodes. To account for this, SEAL (Zhang and Chen, 2018) conditions the message passing on both nodes in the target link by applying a node-labelling trick to the enclosed k-hop neighborhood. They demonstrate that this can result in a suitably expressive GNN for LP. NBFNet (Zhu et al., 2021) conditions the message passing on a single node in the target link by parameterizing the generalized Bellman-Ford algorithm. In practice, it’s been shown that conditional message passing is prohibitively expensive to run on many LP datasets (Chamberlain et al., 2022). Instead, recent methods take the standard GNN representations and an additional pairwise encoding into their scoring function for link prediction. For the pairwise encoding, Neo-GNN (Yun et al., 2021) considers the higher-order overlap between neighborhoods. BUDDY (Chamberlain et al., 2022) estimates subgraph counts via sketching to infer information surrounding a target link. Neural Common-Neighbors with Completion (NCNC) (Wang et al., 2023a) encodes the enclosed 1-hop neighborhood of both nodes. Lastly, LPFormer (Shomer et al., 2024) adapts a transformer to learn the pairwise information between two nodes.
Generalization in Link Prediction: Generalization methods for LP rely on a mix of link and node features in order to improve LP model performance. DropEdge (Rong et al., 2020) randomly removes edges with increasing probability from the training adjacency matrix, allowing for different views of the graph. Edge Proposal Sets (EPS) (Singh et al., 2021) considers two models – a filter and rank model. The filter model is used to augment the graph with top-k relevant common neighbors, while the rank method scores the final prediction. (Wang et al., 2023b) built Topological Concentration (TC), which considers the overlap in subgraph features for a given node with each connected neighbor, correlating well with LP performance for individual links. To improve the performance of links with a low TC, a re-weighting strategy applies more emphasis on links with a lower TC. Counter-Factual Link Prediction (CFLP) (Zhao et al., 2022) conditions a pre-trained model with edges that contain information counter to the original adjacency matrix, allowing models to generalize on information not present in a provided dataset.
3. Benchmark Dataset Construction
(a) CN
(b) PA
(c) SP
Figure 1. An example of the three splitting strategies: (a) Common Neighbors, (b) Preferential-Attachment, (c) Shortest-Path. The dashed lines represent a cut on the source and target node that forms a given edge for our splitting strategy. The color of the line distinguishes the score assigned by the heuristic.
[…]
An example of the three splitting strategies: (a) Common Neighbors, (b) Preferential-Attachment, (c) Shortest-Path. The dashed lines represent a cut on the source and target node that forms a given edge for our splitting strategy. The color of the line distinguishes the score assigned by the heuristic.
In this section, we explain how LPShift induces a shift in each dataset’s structure; clarifying the importance of each structural measure and their application to splitting graph data.
3.1. Types of Distribution Shifts
We induce distribution shifts by splitting the links based on key structural properties affecting link formation and thereby LP. We consider three type of metrics: Local structural information, Global structural information, and Preferential Attachment. Recent work by (Mao et al., 2024b) has shown the importance of local and global structural information for LP. Furthermore, due to the scale-free nature of many real-world graphs and how it relates to link formation (Barabási and Albert, 1999), we also consider Preferential Attachment. A representative metric is then chosen for each of the three types, shown as follows:
(1) Common Neighbors (CNs): CNs measure local structural information by considering only those nodes connected to the target and source nodes. A real-world case for CNs is whether you share mutual friends with a random person, thus determining if they are your “friend-of-a-friend” (Adamic and Adar, 2003). CNs plays a large role in GNN4LP, given that NCNC (Wang et al., 2023a) and EPS (Singh et al., 2021) integrate CNs into their framework and achieve SOTA performance. Furthermore, even on complex real-world datasets, CNs achieves competitive performance against more advanced neural models (Hu et al., 2020). To control for the effect of CNs on shifted performance, the relevant splits will consider thresholds which include more CNs.
(2) Shortest Path (SP): SP captures a graph’s global structural information, thanks to the shortest-path between a given target and source node representing the most efficient path for reaching the target (Russell and Norvig, 2009). The shift in global structure caused by splitting data with SP can induce a scenario where a model must learn how two dissimilar nodes form a link with one another (Evtushenko and Kleinberg, 2021), which is comparable to the real-world scenario where two opponents choose to co-operate with one another (Schelling, 1978; Granovetter, 1978).
(3) Preferential Attachment (PA): PA captures the scale-free property of larger graphs by multiplying the degrees between two given nodes (Barabási and Albert, 1999). When applied to graph generation, PA produces synthetic Barabasi-Albert (BA) graphs which retain the scale-free property to effectively simulate the formation of new links in real-world graphs, such as the World Wide Web (Barabási and Albert, 1999; Albert and Barabási, 2002). Similar to CNs, the relevant PA splits will consider thresholds that integrate higher PA values.
3.2. Dataset Splitting Strategy
In the last subsection we described the different types of metrics to induce distribution shifts for LP. The metrics cover fundamental structural properties that influence the formation of new links. Given the functionality of: CN, SP, and PA within link prediction (Liben-Nowell and Kleinberg, 2003), LPShift applies the three heuristics to induce distribution shifts which respectively target: local, global, and scale-free properties of the graph. Therefore, providing a meaningfully-difficult challenge to GNN4LP, as empirically-demonstrated in Section 5. We now describe how LPShift applies these heuristic measures to split the dataset into train/validation/test splits to induce such shifts.
Figure 2. The Common Neighbors distribution of the HeaRT versus the original testing and validation negative samples from LPShift’s ogbl-collab ’CN - 5,3’ dataset.
[…]
The Common Neighbors distribution of the HeaRT testing and validation negative samples versus the original testing and validation negative samples compared against the positive samples drawn from LPShift’s ogbl-collab ’CN - 5,3’ dataset.
In order to build datasets with structural shift, we apply a given neighborhood heuristic to score each link. This score is then compared to a threshold ($i_{train},i_{valid})$ to categorize a link within a different sample distribution. As denoted in Alg. 1, the heuristic score of the link $(u,v)$ is $h(u,v)$. The link falls into: training when $h(u,v)<i_{train}$, validation when $i_{train}<h(u,v)\leq i_{valid}$, and testing when $h(u,v)>i_{valid}$. The new training graph is constructed from the original OGB dataset (Hu et al., 2020). Validation and testing samples are removed from the new training graph to prevent test-leakage and limited to 100-thousand edges maximum.
Algorithm 1 Dataset Splitting Strategy (LPShift)
1:
2:$G=$ Initial Graph,
[…]
4:$i_{train}$, $i_{valid}=$ Heuristic score thresholds
5:$Train,Valid,Test=\emptyset,\emptyset,\emptyset$
6:while edge, $(u,v)$ not visited in $G$ do
7:     $\Psi(u,v)=h(u,v)$triangleright Score edge with heuristic
8:     if $h(u,v)\leq i_{train}$ then
9:         $Train\leftarrow(u,v)$
10:     else if $h(u,v)>i_{train}\text{ and }h(u,v)\leq i_{valid}$ then
11:         $Valid\leftarrow(u,v)$
12:     $\triangleright$ $h(u,v)>i_{valid}$, Test Split
13:         $Test\leftarrow(u,v)$
14:     end if
15:end while
16:return $Train,Valid,Test$triangleright Return Final Splits
With Figure 1, we provide a small example of how splits are produced by our proposed splitting strategy. Specifically, Figure 1(a) demonstrates an outcome of the CN split labelled “CN - 1,2” where sampled edges pulled from the: black-dotted line = training (no CNs), red-dotted line = validation, (1 CN), and blue-dotted line = testing ($\geq$2 CNs). Figure 1(b) follows a similar high-level process, focusing on node-degree instead of CNs. As such, the validation (red) cut splits nodes with degree 3 and the testing (blue) cut splits nodes with degree 1. Figure 1(c) details Shortest-Path (SP), where the red cut works on a path of length 3 and is categorized by taking SP as the denominator of the ratio, $\frac{1}{SP}$.
To test how different LPShift thresholds impact performance, we adjust the $i_{train}$ and $i_{valid}$ thresholds to produce 3 varied CN and PA splits; as well as 2 varied SP splits. The variations in split thresholds were chosen based on two conditions. 1) Structural information within splits varies due to user-defined thresholds. In the ”Forward” scenario (i.e. CN - 1,2), splits are given increasingly more structural information between training to validation, giving the model an easier time generalizing under testing. For the ”Backward” scenario (i.e, CN - 2,1), stricter thresholds mean less structural information between validation to testing; making generalization difficult. 2) The final dataset split contains a sufficient number of samples. Each LPShift split requires enough split samples to allow model generalization. Given the limited number of SP split samples, the SP splits were limited to 2 variants.
3.3. Why use HeaRT with LPShift?
It is common practice for most link prediction baselines to randomly-sample all negatives samples. However, random sampling is shown to be unrealistic, given that it is trivially-easy for a GNN4LP model trained via supervised learning to distinguish positive and negative samples based solely on the sample’s structure (Li et al., 2024). We demonstrate an example of this effect in the right subplot of Figure 2, where the original setting samples negative at random to overwhelmingly contain 0 Common Neighbors. The samples produced by the original setting can be especially misleading, GNN4LP models could achieve high performance due to this overwhelming density of negative samples with 0 CNs versus positive samples with $>$0 CNs. Figure 2’s left subplot shows how HeaRT corrects this problem by sampling valid and test negatives whose distribution of CNs overlap with their respective positive samples (shown in red and blue), as well as the distribution of positive training samples (shown in black on both subplots). Therefore, HeaRT sampling avoids a critical pitfall within the original setting, where GNN4LP models could learn the trivial dynamic where any sample with more than 0 CNs as a positive edge and any sample with 0 CNs as a negative edge.
4. Why LPShift? A Theoretical Perspective
Section 3 details how heuristic choice and HeaRT influence LPShift’s ability to alter dataset structure. We now consider how GNNs, as neural backbones for GNN4LP, rely on assumptions about the dataset’s underlying sample distributions to learn representations for effective link prediction. From these assumptions, GNN4LP capture structural heuristics associated with link formation, such as LPFormer estimating Katz-Index (Shomer et al., 2024). LPShift undermines these assumptions through direct manipulation of the dataset sample distribution, which is quantifiable within probability theory. Under which, we treat a link predictor as a function to learn a probability measure over a pairwise embedding. This learned probability measure, coupled with the obtained embedding, determines the ability of GNN4LP to quantify whether link formation occurs or not. We provide a theoretical perspective on this effect below:
Notation:
We denote the node feature space as $\textbf{X}\in\mathbb{R}^{N\times F}$, where $F$ is the number of node features, and the adjacency matrix representing a given graph dataset as, $\textbf{A}\in\{0,1\}^{N\times N}$.
Definition 4.1.
[…]
Theorem 4.2.
[…]
Proof.
[…]
Indeed, when the i.i.d assumption across dataset samples holds, supervised training of neural LP models with GNN backbones enables message-passing across training samples. This further allows GNN4LP to observe and subsequently model link formation (Bronstein et al., 2021; Zečević et al., 2021), contributing to high downstream performance. However, Theorem 4.2 conditions link formation dynamics within a given graph dataset, meaning that the necessary i.i.d. subsets for learning and then achieving generalized performance are obscured. In effect, the GNN4LP model gets a limited view on pairwise information contained within the dataset. As such, LPShift allows direct testing of the hypothesis: Can GNN4LP model structural heuristics for effective downstream performance in scenarios with limited structural information? We verify this effect empirically in Section 5.
5. Experiments
To bridge the gap for GNN4LP generalizing under distribution shifts, this work addresses the following questions: (RQ1) Is the distribution shift induced by LPShift significant? (RQ2) Can SOTA GNN4LPs generalize under our proposed distribution shifts? (RQ3) Can current generalization methods boost the performance of GNN models? (RQ4) What components of the proposed distribution shift are affecting the LP model’s performance?
5.1. Experimental Setup
Datasets: We consider 16 “Forward” and “Backward” LPShift splits for the following 3 OGB datasets (Hu et al., 2020): ppa, ddi, and collab, for a total of 48 tested splits. The resulting datasets represent tasks in three separate domains and three shifted scenarios, allowing a comprehensive study of LP generalization under distribution shift. For all datasets, we create multiple splits corresponding to each structural property, as detailed in Section 3.2. For the “Forward” split, denoted as $(X,Y,Z)$, an increase in $Y$ and $Z$ indicates more structural information available to the training adjacency matrix. The “Backward” split swaps the training and testing splits from their counterpart in the “Forward” split, resulting in the training adjacency matrix losing access to structural information as $X$ and $Y$ increase.
GNN4LP Methods: We test multiple SOTA GNN4LP methods including: NCNC (Wang et al., 2023a), BUDDY (Chamberlain et al., 2022), LPFormer (Shomer et al., 2024), SEAL (Zhang et al., 2021) and Neo-GNN (Yun et al., 2021). We further consider GCN (Kipf and Welling, 2017) as a simpler GNN baseline, along with the Resource Allocation (RA) (Zhou et al., 2009) heuristic. All models were selected based on their benchmark performance with the original OGB datsets (Hu et al., 2020) and their architectural differences detailed in Section 2.
Graph-Specific Generalization Methods: We also test the performance of BUDDY with multiple generalization techniques. This includes DropEdge (Rong et al., 2020), which randomly removes a portion of edges from the training adjacency matrix. Edge Proposal Sets (EPS) (Singh et al., 2021), which uses paired LP models to filter and rank top-k edges to insert as Common Neighbors in the training adjacency matrix. Lastly, we consider Topological Concentration (TC) (Wang et al., 2023b), which re-weights the edges within the training adjacency matrix based on their shared information. Note: We were unable to test CFLP (Zhao et al., 2022) and EERM (Wu et al., 2022) after adapting their current implementations to the HeaRT evaluation setting. CFLP experienced an out-of-memory error on all tested LPShift splits. EERM experienced an out-of-memory error on every LPShift split of ogbl-ppa and ogbl-ddi, while also exceeding 48 hours per run on ogbl-collab without converging.
Traditional Generalization Methods: The final round of tests applies: IRM (Arjovsky et al., 2019), VREx (Krueger et al., 2021), GroupDRO (Sagawa et al., 2019), DANN (Ganin et al., 2016), and DeepCORAL (Sun and Saenko, 2016) as generalization methods to GCN. Given that LPShift splits graph data directly on the sample structure, all tests with traditional generalization methods treat training samples as an environmental partition. The partitions =$\mathcal{E}=\{e_{1},...,e_{N}\}$ are composed of $N$ environmental subsets, where each subset is binned based on their neighborhood heuristic; as demonstrated for CNs in Figure 3. This binning of environmental subsets forces traditional generalization methods to target graph structure, controlling for the method’s ability to improve GNN4LP performance under structural shift induced by LPShift.
Figure 3. The original environment of dataset samples (shown left) split into subsets by their CN heuristic scores (shown right).
[…]
The original environment of dataset samples (shown left) split into subsets by their CN heuristic scores (shown right).
Evaluation Setting: We consider the standard evaluation procedure in LP, in which every positive validation/test sample is compared against $M$ negative samples. The goal is that the model should output a higher score (i.e., probability) for positive sample than the negatives. To create the negatives, we make use of the HeaRT evaluation setting (Li et al., 2024) which generates $M$ negatives samples per positive sample according to a set of common LP heuristics. In our study, we set $M=250$ and use CNs coupled with PPR as the heuristic in HeaRT.
Evaluation Metrics: We evaluate all methods using the mean reciprocal rank (MRR) across multiple Hits@K ranking metrics, K=$\{1,3,5,10,20,50,100\}$.
Table 1. The average percent change in MRR for each tuned model versus the HeaRT benchmark (Li et al., 2024), sorted by dataset then LPShift split and direction.
Dataset Split Direction GCN RA NCNC NeoGNN BUDDY LPFormer SEAL
PPA CN Forward -69.4% -84.0% -84.5% -77.0% -84.8% -91.2% -67.5%
Backward -91.4% -96.9% -75.0% -96.3% -92.0% -88.3% -96.3%
SP Forward -78.9% -7.5% -82.2% -72.4% -83.3% -81.7% -75.1%
Backward -75.4% -97.9% -89.8% -86.5% -76.6% -91.4% -96.6%
PA Forward -84.3% -88.5% -85.4% -35.0% -84.6% -77.3% -87.4%
Backward -89.6% -78.5% -74.0% -77.8% -90.3% -75.9% -86.8%
DDI CN Forward -66.1% -50.5% ¿24hr -85.3% -82.4% -93.1% -89.9%
Backward -91.2% -93.9% ¿24hr -80.7% -91.2% -73.4% -93.0%
PA Forward -84.0% -56.2% ¿24hr -86.0% -85.3% -97.1% -87.8%
Backward -80.5% -60.8% ¿24hr -75.1% -83.9% -74.4% -74.4%
Collab CN Forward -5.7% +8.8% -41.5% -49.8% -49.1% +89.6% -29.8%
Backward -69.6% -74.8% -64.1% -79.6% -87.0% -37.7% -81.6%
SP Forward -23.0% +20.4% -49.0% -72.7% -49.7% +33.2% -90.7%
Backward -7.7% -97.7% -95.6% -83.8% -84.0% -67.1% -96.2%
PA Forward -17.5% +3.9% -27.1% -35.7% -50.3% +117.5% -2.4%
Backward -6.5% +41.6% +1.9% -77.1% -44.7% +203.4% +52.9%
Hyperparameters: All methods were tuned on permutations of learning rates in $\{1e^{-2},1e^{-3}\}$ and dropout in $\{0.1,0.3\}$. Each model was trained and tested over five seeds to obtain the mean and standard deviations of their results. Given the significant time complexity of training and testing on LPShift variants for the ogbl-ppa datasets, NCNC and LPFormer were tuned on a single seed, followed by evaluation of the tuned model on five separate seeds. Additional hyperparameter tuning details are included in Appendix A.
5.2. Results for GNN4LP
In order to provide a unified perspective on how distribution shift affects link prediction models, GNN4LP models were trained and tested across five seeded runs on versions of ogbl-collab, ppa, and ddi split by: Common Neighbors, Shortest-Path, and Preferential-Attachment. Examining the results, we have the following three key observations.
Observation 1: Drastic Changes in Model Performance. As shown in Table 1, the majority of synthetic splits generated by LPShift reduce model performance significantly below the original HeaRT standard (Li et al., 2024). This is especially pronounced in denser graphs like ogbl-ppa and ogbl-ddi, indicating that LPShift affects the varied structures of graph data, which then reduces downstream perfomance. The only exception is ogbl-collab, where: RA, NCNC, LPFormer, and SEAL experience measurable improvements over the HeaRT baseline. Given ogbl-collab’s lack of density relative to ogbl-ppa and ogbl-ddi, there are considerations about how LPShift induces a covariate shift to features (Koh et al., 2021), especially where structure is correlated to the feature distribution.
Additional analysis shown in Figure 6 indicates how dataset structure relates to the performance effects of LPShift.
Observation 2: Performance Differs Across LPShift Splits. As shown in Figure 4, regardless of whether a model is tested on a “Forward” or “Backward” split of ogbl-ppa; the change in structural information for each subsequent split gradually changes a model’s performance. On the “CN” split, a stark increase is seen between ”Split 1” and ”Split 2” (also detailed in descending order within Table 10), indicating that more structural information between training and validation improves GNN4LP performance (Wang et al., 2023a), even on dense datasets like ogbl-ppa. The fact that these results include splits produced via Preferential-Attachment, Global Structural Information (SP), and Local Structural Information (CN) indicates the effect of any change in structural information when training LP models (Mao et al., 2024b). In principle, Observation 1 and 2 answer RQ1 and RQ2 by indicating that structural alignment between positive samples from separate splits is critical for GNN4LP performance. LPShift makes this effect quantifiable through user-controlled thresholds, enabling future research to: induce, diagnose, and correct problems caused by LP-specific distribution shift.
Figure 4. The mean scores of the best-performing GNN4LP models on LPShift’s ogbl-ppa dataset. Each line represents a given dataset and split, arranged uniformly between figures. In the case of decreasing performance, the model with the highest average values was selected.
[…]
The mean scores of the best-performing GNN4LP models tested with our proposed splitting strategy. Each line represents a given dataset and split, arranged uniformly between figures. In the case of decreasing performance, the model with the highest average values was selected.
Observation 3: Performance Impact Varies by Model. All raw MRR scores are stored in Appendix F within  Tables 9 to 11. From those tables, we can make the following observations by split type. Common Neighbors: Most models fail to generalize on the “Backward” CN splits. However, once more Common Neighbors are made available in the CN - 4,2 and CN - 5,3 splits; NCNC performs 2 to 3 times better than other GNN4LP models. Therefore, indicating that it is possible to generalize with limited local information. Shortest-Path: GNN4LP Models which rely more on local structural information (i.e. NCNC, LPFormer, and SEAL) typically suffer more under the “Backward” SP splits, resulting in the models performing 2x to 4x worse than BUDDY or GCN. Therefore, indicating the necessity for models to adapt in scenarios with an absence of local structural information. Preferential-Attachment: LPFormer performance on the PA split is 2 times higher than HeaRT’s ogbl-collab (Li et al., 2024), but reduces drastically on LPShift’s ogbl-ppa or ogbl-ddi. Therefore, indicating the impact that structural shift incurs with the increasing density of datasets. These observations within Section 5.2 provide implications for RQ2, showing that targeted model updates; like NCNC applying completion with CNs (Wang et al., 2023a) or LPFormer’s PPR-aware attention (Shomer et al., 2024; Brin and Page, 1998) aid generalization within CN and PA distribution shifts.
5.3. Results for Generalization Methods
In this section, we apply DropEdge (Rong et al., 2020), EPS (Singh et al., 2021), and TC (Zhao et al., 2023) on the previously benchmarked BUDDY (Chamberlain et al., 2022) to determine the feasibility of improving the LP models’ generalization under LPShift.
Table 2. The raw MRR change for the BUDDY results in Table 1, after applying graph-specific generalization methods, averaged over LPShift splits.
Forward Backward
CN SP PA CN SP PA
Collab
ED -0.5 -1.6 -0.11 -0.72 -1.09 +0.36
TC -3.63 -8.02 -2.26 +1.29 +1.34 -7.38
EPS -4.28 -7.82 -4.46 +1.38 +1.45 -8.55
PPA
ED -0.02 +1.31 +0.003 +0.003 0.0 +0.006
TC -1.31 +0.86 -0.90 +1.27 +1.58 -1.09
EPS -1.12 +1.14 -1.02 +1.06 +2.2 -0.91
DDI
ED 0.0 0.0 0.0 +3.87
TC -1.17 0.0 -0.12 -0.46
EPS -0.86 -0.58 +1.19 +7.5
Observation 1: Graph-Specific Generalization Methods Can Help. As demonstrated in Table 2, the two generalization methods specific to LP: TC (Wang et al., 2023b) and EPS (Singh et al., 2021) fail to increase performance under LPShift in the ”Forward” CN and PA splits. However, the methods do improve performance for the ”Backward” DDI split, where both EdgeDrop and EPS increase model performance by 3.87% and 7.5%, respectively. This is likely due to the exceptionally dense nature of DDI, where numerous high-degree nodes are likely made more relevant to performance through: 1) the removal of edges from EdgeDrop or 2) connecting relevant neighbors with EPS. To validate this, we calculate Earth Mover’s Distance (EMD) (Rubner et al., 1998) between the heuristic scores of the training and testing splits before and after applying the generalization methods. EPS injects CNs into the training adjacency matrix, significantly altering the training and testing distributions. This drastic change is indicated in Figures 5 with the 28-92% improvement of EMD scores between training and testing from applying EPS. Such a change in EMD, coupled with the Table 2 results, demonstrates that generalizing under LPShift can be achieved with targeted augmentations to the graph data structure. However, EdgeDrop has relatively-minimal effect on the EMD scores, indicating that non-targeted augmentations can generalize from highly-dense training subsets to less-dense test subsets.
TC often decreases performance on ”Forward” splits. This is likely due to LPShift’s distinct split thresholds; meaning there is limited structural overlap between sample distributions. As such, TC can’t re-weight the training adjacency matrix for improved generalization to neighborhood information (Wang et al., 2023b; Li et al., 2022a). This result runs contrary to current work, where re-weighting is effective for handling distribution shifts in other graph tasks (Zhou et al., 2022b) and computer vision (Fang et al., 2020). However, the ”Backward” CN and SP splits indicate consistent performance increases induced by TC.
Figure 5. The EMD values calculated between the heuristic scores of training and testing samples. Note: The tested heuristics correspond to their labelled splits, so as to simulate the dataset splitting.
[…]
The EMD values calculated between the heuristic scores of training and testing samples. Note: The tested heuristics correspond to their labelled splits, so as to simulate the dataset splitting.
Observation 2: Traditional Generalization Methods Work in Limited Scenarios. As shown in Table 3, any of the generalization methods tested on the ”Forward’ split rarely improve performance. IRM and DRO often reduce performance the most, sometimes by more than 10% MRR. All methods encounter more consistent performance improvements on the ”Backward” ogbl-ppa and ogbl-ddi splits. Given that all generalization methods are trained on environmental subsets defined by structural heuristics, it follows that the traditional generalization methods improve performance under LPShift when given sample subsets with distinct and diverse structural patterns. Future work could explore this idea further, applying different heuristic scores to each environmental subset or extending invariant learning principles to work for the link prediction task (Wu et al., 2024). As shown in Table 3, IRM performs well on ”Backward” ogbl-ddi splits but fails on ”Backward” ogbl-collab splits. Indeed, there is no single generalization method which functions under all LPShift scenarios for each given dataset. This result is consistent with the application of OOD generalization techniques in other data domains (Gulrajani and Lopez-Paz, 2020). As such, this result also answers RQ3 by showing LPShift poses a challenge to current generalization methods.
Table 3. The raw MRR change for the GCN results in Table 1, after applying traditional generalization methods, averaged over LPShift splits.
Forward Backward
CN SP PA CN SP PA
Collab
IRM -10.62 -8.51 -6.74 -2.66 -2.18 -14.86
VREx -0.53 +0.39 +2.12 +0.31 +0.02 -1.29
DRO -11.96 -7.76 -7.41 -2.51 -1.32 -13.66
DANN -0.13 +0.74 +2.35 +0.39 +3.36 -1.75
CORAL -0.54 +0.36 +2.02 +0.25 +3.32 -1.29
PPA
IRM -6.16 +0.02 -0.15 +0.3 -4.51 +0.53
VREx -0.9 +0.08 +0.07 +0.79 -0.17 +0.03
DRO -5.86 -1.87 -0.73 +0.45 -4.48 +0.83
DANN -0.84 +0.13 +0.65 +0.72 -0.12 -0.45
CORAL -0.91 +0.02 +0.37 +0.71 -0.17 -0.51
DDI
IRM -3.44 -0.56 +3.24 +3.26
VREx -3.68 -0.66 +2.76 +1.56
DRO -3.47 -0.77 +3.16 +1.89
DANN -3.55 -0.66 +3.26 +1.59
CORAL -3.72 -0.62 +2.74 +1.18
5.4. Discussion
Does GNN4LP generalize and do generalization methods work? As detailed in Table 1, LPShift significantly reduces the performance for a majority of GNN4LP models versus the HeaRT baseline. This is especially notable given the difficulty HeaRT imposes on the original benchmark setting (Hu et al., 2020). This observation, coupled with the limited success of all generalization methods on ”Forward” splits in Tables 2 and 3, shows the difficulty posed on LP models when trained with limited structural information relative to evaluation samples.
How is the proposed distribution shift affecting performance? In order to answer RQ4 and verify that LPShift’s effect on dataset structure is significantly correlated to performance, Figure 6 follows a similar analysis conducted in (Wang et al., 2023a). In which, CN’s predictive performance is measured and compared under the ”Forward” and ”Backward” LPShift. We choose CN as the predictor, given that directly ranks given dataset samples by structure to predict if a link forms; meaning it’s predictive performance is sensitive to a change in dataset structure.
Figure 6. Three subplots corresponding to: Top.) CN predictor performance for the original ogbl-collab with it’s respective LPShift ’CN - 1,2’, and ’CN - 2,1’ splits., Bottom-Left.) the density of Common Neighbors for ’CN - 1,2’ LPShift on ogbl-collab, Bottom-Right.) the density of Common Neighbors for ’CN - 2,1’ LPShift on ogbl-collab.
[…]
The bottom two subplots in Figure 6 depict the density estimates for CN distributions of LPShift’s CN - 1,2 and CN - 2,1 splits on the ogbl-collab dataset. The topmost subplot depicts the performance of CN on the datasets depicted in the bottom two subplots, as well as the original ogbl-collab dataset.
Figure 6 demonstrates the targeted effect of LPShift on dataset structure, the reduced performance (1% MRR) of the CN ranking on the CN - 1,2 split is due to the lack of structural information in the training split (CN = 0), limiting CN’s ability to rank valid and test samples with $>$1 CNs. The structure-to-performance link is clearer in the CN - 2,1 scenario: where training with 2+ CNs yields 60% MRR, but fails to generalize on test samples with zero CNs, shown by the 1% MRR score. Given that all GNN4LP models incorporate neighborhood/structural information into their architectures (Wang et al., 2023a; Chamberlain et al., 2022), it follows that GNN4LP’s reduced performance under LPShift is due to the splitting strategy enforcing a scenario where models must generalize to dataset samples drawn from a different distributions, empirically verifying Theorem 4.2.
6. Conclusion
This work proposes LPShift, a simple dataset splitting strategy for inducing structural shift relevant for link prediction. The effect of this structural shift was then benchmarked on 16 shifted versions of ogbl-collab, ppa, and ddi, posing a unique challenge for SOTA GNN4LP models and generalization methods versus the HeaRT baseline. Further analysis indicates that current generalization methods improve performance under LPShift in scenarios with diverse pairwise structures amongst training samples. This indicates promising future directions for targeted graph data augmentation and invariant learning methods to improve generalization under distribution shift in the link prediction task. Regarding this, LPShift also demonstrates that GNN4LP as standalone models are vulnerable to distribution shift induced by pairwise heuristics. Given the severity of structural shift induced by LPShift, future work can consider relaxing the threshold between sample distributions or consider mixing structural heuristics to better mimic real-world distribution shift and the complexities therein.
Acknowledgements.
All authors are supported by the National Science Foundation (NSF) under grant numbers DGE2244164, CNS2321416, IIS2212032, IIS2212144, IOS2107215, DUE2234015, CNS2246050, DRL2405483 and IOS2035472, US Department of Commerce, Gates Foundation, the Michigan Department of Agriculture and Rural Development, Amazon, Meta, and SNAP.
Appendix A Additional Training Details
This section provides relevant details about training and reproducing results not mentioned in Section 5.1:
  • Please consult the project README for building the project, loading data, and re-creating results.
  • All models and link predictors are fixed to 3 layers in their neural architectures.
  • GCN and LPFormer were tuned to 128 hidden dimensions.
  • BUDDY, NCNC, NeoGNN, and SEAL were tuned to 256 hidden dimensions.
  • All experiments were conducted with a single A6000 48GB GPU and 1TB of available system RAM.
  • We apply the ’NCNC2’ variant of NCNC with an added depth argument of 2 (Wang et al., 2023a) for all CN and SP splits of ogbl-collab.
  • Otherwise, we set depth equal to 1, so as to reduce runtime.
  • NeoGNN use 1-hop neighborhoods on all tested datasets.
  • The learning rate and dropout for initial batch size tuning was fixed at $1e^{-3}$ and $0.1$ respectively.
[…]
  • PPA Training Edges are limited to 3 million edges total.
  • Validation and testing edges that are duplicated with training edges are removed from the edge index.
  • In order to provide overlap within a given dataset, validation and testing edges that do not connect to training nodes are removed from the edge index.
  • After sampling the necessary training edges, the adjacency matrix is extracted from the edge index, converted to an undirected graph and has any edge weights standardized to 1.
Common Neighbors, Preferential-Attachment and Shortest-Path, as respectively shown in Figure 1, are interchangeable within the dataset splitting strategy. Details about how Common Neighbors functions within the strategy are included in Section 3.2.
As another possible component of LPShift, Preferential-Attachment determines the degrees between a given source and target node and then multiples the two to produce the score, based on that score, the sample is then sorted into a new dataset split.
As another possible component of LPShift, Shortest-Path determines the score by determining the minimum number of nodes necessary to reach the target node from the source node. If there is a link between the two nodes, we remove the link and then re-add to the adjacency matrix after the score calculation. The final Shortest-Path score applies the calculated shortest-path length, $SP(u,v)$ as the denominator in a ratio of $\frac{1}{SP(u,v)}$, which is then used to sort the sample into it’s respective dataset split.
Appendix B Cosine Similarity of ogbl-collab
Given LPShift’s effect on structure, we also quantify the extent to which it can affect the feature distributions of the real-world features within the ogbl-collab data. Note: ogbl-ddi was not considered due to a lack of node features and ogbl-ppa consists of one-hot encoded features derived directly from structure.
Table 4. The cosine-similarity table of ogbl-collab’s original feature distribution and its LPShift versions.
Split Train/Test Train/Valid Valid/Test
Original 83.50 ± 7.33 83.40 ± 7.33 86.91 ± 6.58
CN (0,1,2) 87.73 ± 6.12 87.42 ± 6.17 91.78 ± 4.01
(0,2,4) 87.06 ± 6.08 86.72 ± 6.14 91.01 ± 4.29
(0,3,5) 86.45 ± 6.12 86.19 ± 6.12 90.42 ± 4.41
(2,1,0) 85.53 ± 6.74 85.22 ± 6.79 90.45 ± 4.49
(4,2,0) 85.97 ± 7.09 85.55 ± 7.15 90.71 ± 4.43
(5,3,0) 86.08 ± 7.29 85.62 ± 7.34 90.74 ± 4.53
SP ($\infty$,6,4) 84.11 ± 6.94 82.12 ± 7.34 85.60 ± 6.75
($\infty$,4,3) 84.35 ± 6.93 82.66 ± 7.45 85.97 ± 6.75
(4,6$\infty$) 81.10 ± 7.74 82.15 ± 7.79 81.86 ± 7.48
(3,4$\infty$) 81.99 ± 7.84 82.81 ± 7.98 82.76 ± 7.72
PA (0,50,100) 83.92 ± 7.22 82.59 ± 7.63 85.50 ± 7.14
(0,100,200) 87.48 ± 6.16 87.55 ± 6.10 92.52 ± 3.70
(0,150,250) 87.61 ± 6.10 87.81 ± 6.17 92.45 ± 3.84
(100,50,0) 82.28 ± 7.96 83.22 ± 8.00 83.17 ± 7.85
(200,100,0) 84.75 ± 6.65 85.32 ± 6.64 89.94 ± 4.52
(250,150,0) 85.08 ± 6.65 85.51 ± 6.64 90.62 ± 4.46
Appendix C Time to Split Tested Dataset Samples
A key consideration for LPShift’s application as a splitting strategy is to alleviate the burden of gathering new datasets, allowing researchers to control for and then induce a distribution shift in link-prediction datasets quickly; without requiring an expensive and time-consuming project to build a new dataset. This consideration is inspired by current graph and node-classification benchmark datasets, all of which induce distribution shifts in pre-existing benchmark datasets (Gui et al., 2022),(Ji et al., 2022),(Koh et al., 2021). LPShift is not meant to replace high-quality benchmark datasets, especially for distribution shifts, but to serve as a supplement for current datasets and enhance understanding of LP generalization. Results demonstrating LPShift’s time-efficiency on tested dataset splits are included below in Table 5.
Table 5. The average time in seconds (s) across 10 runs to generate each ’Forward’ and ’Backward’ split for ogbl-ppa, ogbl-collab, ogbl-ddi.
Split ogbl-collab ogbl-ppa ogbl-ddi
CN (0, 1, 2) 7.48 s 177.89 s 13.45 s
(0, 2, 4) 7.49 s 177.09 s 14.21 s
(0, 3, 5) 7.63 s 178.23 s 13.64 s
(2, 1, 0) 7.66 s 184.98 s 13.31 s
(4, 2, 0) 7.63 s 186.49 s 13.71 s
(5, 3, 0) 7.86 s 185.95 s 13.42 s
SP ($\infty$, 6, 4) 53.12 s 2748.64 s
($\infty$, 4, 3) 52.24 s 2705.91 s
(4, 6, $\infty$) 53.93 s 2715.34 s
(3, 4, $\infty$) 53.7 s 2751.56 s
PA (0, 50, 100) 19.25 s 406.5 s 15.41 s
(0, 100, 200) 19.05 s 408.04 s 14.55 s
(0, 150, 250) 19.35 s 407.81 s 13.96 s
(100, 50, 0) 19.65 s 425.3 s 14.42 s
(200, 100, 0) 19.34 s 409.93 s 14.59 s
(250, 150, 0) 19.37 s 403.55 s 14.05 s
Appendix D Size of Dataset Samples
In order to aid reproducibility of results when running LPShift manually; this section details the number of training, validation, and test edges for all LPShift dataset split types introduced within in Section 3 and all datasets.
Table 6. The number of ogbl-collab samples for the forward and backward heuristic splits.
CN (0,1,2) (0,2,4) (0,3,5) (2,1,0) (4,2,0) (5,3,0) Train 57638 237928 493790 1697336 1193456 985820 Valid 6920 20045 31676 23669 24097 25261 Test 4326 14143 21555 9048 11551 11760 SP ($\infty$,6,4) ($\infty$,4,3) (4,6$\infty$) (3,4$\infty$) Train 46880 52872 1882392 1877626 Valid 1026 1238 5222 4384 Test 2759 3457 2626 7828 PA (0,50,100) (0,100,200) (0,150,250) (100,50,0) (200,100,0) (250,150,0) Train 210465 329383 409729 1882392 1877626 457372 Valid 46626 62868 75980 64729 64202 65323 Test 9492 25527 39429 41381 30983 30999
Table 7. The number of ogbl-ppa samples for the forward and backward heuristic splits.
CN (0,1,2) (0,2,4) (0,3,5) (2,1,0) (4,2,0) (5,3,0) Train 2325936 3000000 3000000 3000000 3000000 3000000 Valid 87880 95679 98081 96765 93210 92403 Test 67176 83198 88778 92798 85448 81887 SP ($\infty$,6,4) ($\infty$,4,3) (4$\infty$,6) (3,4$\infty$) Train 17464 134728 3000000 3000000 Valid 149 4196 90511 97121 Test 20 1180 458 74068 PA (0,5k,10k) (0,10k,20k) (0,15k,25k) (10k,5k,0) (20k,10k,0) (25k,15k,0) Train 95671 98562 99352 3000000 3000000 3000000 Valid 95671 98562 99352 90623 89671 91995 Test 45251 63178 72382 44593 34321 35088
Table 8. The number of ogbl-ddi samples for the forward and backward heuristic splits.
CN (0,1,2) (0,2,4) (0,3,5) (2,1,0) (4,2,0) (5,3,0) Train 3938 5636 6580 10000 10000 10000 Valid 1184 992 942 966 1668 1462 Test 212 525 809 1808 2296 2832 PA (0,5k,10k) (0,10k,20k) (0,15k,25k) (10k,5k,0) (20k,10k,0) (25k,15k,0) Train 5132 8246 11718 10000 10000 10000 Valid 1572 3208 5058 2678 6188 6300 Test 419 1422 3694 1900 3632 5970
Appendix E Is LPShift effective at inducing distribution shift?
The following section will explore the capability of LPShift to induce a measurably-significant distribution shift in the structure of the ogbl-collab dataset. We apply the 2-sample Kolmgorov-Smirnov (KS) test (Hodges Jr, 1958) to compare if training, validation, and testing distributions can be sampled from one another, both before and after applying LPShift. As a controlled baseline to test LPShift, we generate a Holme-Kim (HK) graph (Holme and Kim, 2002) with a 40% chance to close a triangle, allowing the HK graph to contain numerous Common Neighbors without becoming fully-connected. HK graph generation parameters are included in Appendix B.
The Holme-Kim graph used for analysis in Figure 7 was generated with the following parameters:
  • $n$ = $235868$, $m$ = $5$, $p_{c}$ = $0.4$, seed = $42$
Figure 7. Three subplots detailing CN distributions for: 1.) the unaltered ogbl-collab dataset 2.) a Holme-Kim (HK) graph with a random split 3.) the HK graph from 2. split with LPShift’s CN - 5,3 strategy
[…]
Three subplots detailing CN distributions for: 1.) the unaltered ogbl-collab dataset 2.) a Holme-Kim (HK) graph with a random split 3.) the HK graph from 2. split with LPShift’s CN - 5,3 strategy
The first subplot in Figure 7 extends the reasoning introduced with NCNC (Wang et al., 2023a). As such, this subplot indicates there is a natural shift for CNs within the original ogbl-collab dataset (Hu et al., 2020). The p-value of 0, measured across both split permutations, indicates that the training distribution of CNs is shifted from the validation and testing distributions. The second subplot depicts a randomly-split HK graph (Holme and Kim, 2002), where CN distributions for each split match one another, further indicated by the p-values of 1. The third subplot depicts the HK graph from the second subplot split with LPShift’s CN - 5,3 strategy, resulting in a distinct shift between all dataset splits, as confirmed by the 0 p-values. As such, LPShift induces structural shift that is as measurably dissimilar as the structural shift present in the original ogbl-collab dataset, even when the initial dataset splits are measurably identical. Additionally, the CN - 5,3 split causes the shape of the HK graph’s CN distributions to become more similar to the shift observed within the original ogbl-collab dataset, indicating that the ”Backward” LPShift strategy can function like a real-world distribution shift.
Appendix F Dataset Results
This section includes Tables 9 to 11, displaying baseline results for initial LPShift experiments.
Table 9. ogbl-collab results reported in MRR with the best bolded and the second best underlined.
Split Models RA GCN BUDDY NCNC LPFormer NeoGNN SEAL CN (0, 1, 2) 32.22 14.89 ± 0.67 17.48 ± 1.19 5.34 ± 2.54 4.27 ± 1.17 5.53 ± 1.52 7.06 ± 1.67 (0, 2, 4) 29.74 17.75 ± 0.44 15.47 ± 0.57 13.99 ± 1.35 12.97 ± 1.24 13.26 ± 0.39 20.60 ± 6.23 (0, 3, 5) 29.86 19.08 ± 0.26 16.60 ± 0.89 15.06 ± 1.43 25.36 ± 2.04 14.21 ± 0.88 19.78 ± 1.24 (2, 1, 0) 0.6 5.04 ± 0.13 3.70 ± 0.13 1.46 ± 0.02 2.01 ± 0.95 2.16 ± 0.05 1.01 ± 0.02 (4, 2, 0) 4.79 4.87 ± 0.17 3.55 ± 0.09 6.13 ± 0.49 3.87 ± 0.74 3.41 ± 0.16 3.38 ± 0.62 (5, 3, 0) 15.9 6.78 ± 0.19 5.40 ± 0.04 12.87 ± 1.50 8.36 ± 0.63 7.85 ± 0.74 8.08 ± 1.96 SP ($\infty$, 6, 4) 33.87 12.64 ± 0.85 16.20 ± 1.40 14.43 ± 1.36 4.6 ± 3.15 6.14 ± 1.44 2.08 ± 1.50 ($\infty$, 4, 3) 33.91 15.50 ± 0.16 16.42 ± 2.30 18.33 ± 1.24 15.7 ± 2.87 5.82 ± 1.20 2.10 ± 1.24 (4, 6, $\infty$) 0.69 4.75 ± 0.10 6.73 ± 0.32 0.73 ± 0.13 3.16 ± 0.62 4.55 ± 0.27 0.93 ± 0.07 (3, 4, $\infty$) 0.63 3.81 ± 0.16 3.71 ± 0.39 0.95 ± 0.24 1.86 ± 0.48 2.54 ± 0.09 0.80 ± 0.01 PA (0, 50, 100) 36.87 19.81 ± 0.59 21.27 ± 0.74 17.24 ± 0.17 25.31 ± 5.67 15.94 ± 0.49 29.06 ± 1.57 (0, 100, 200) 26.78 12.68 ± 0.20 14.04 ± 0.76 12.34 ± 3.01 11.98 ± 3.12 13.50 ± 0.84 20.69 ± 1.51 (0, 150, 250) 24.07 12.73 ± 0.32 13.06 ± 0.53 12.01 ± 0.82 12.43 ± 6.62 12.81 ± 0.51 16.23 ± 3.69 (100, 50, 0) 33.09 24.33 ± 0.537 24.95 ± 0.92 20.81 ± 1.86 17.76 ± 2.01 6.42 ± 0.462 31.83 ± 6.44 (200, 100, 0) 42.28 13.97 ± 0.52 15.52 ± 0.73 18.62 ± 1.54 27.56 ± 9.10 4.20 ± 0.37 31.96 ± 6.65 (250, 150, 0) 44.14 12.98 ± 0.45 13.36 ± 0.83 18.71 ± 0.99 24.04 ± 11.35 4.40 ± 0.63 39.58 ± 4.81
Table 10. ogbl-ppa results reported in MRR with the best bolded and the second best underlined.
Split Models RA GCN BUDDY NCNC LPFormer NeoGNN SEAL CN (0,1,2) 4.71 7.81 ± 0.08 7.90 ± 0.32 7.08 ± 3.08 3.28 ± 0.63 3.49 ± 0.10 11.91 ± 1.85 (0,2,4) 4.45 8.18 ± 0.12 3.83 ± 0.24 9.89 ± 1.92 2.46 ± 0.51 5.79 ± 0.42 4.84 ± 0.10 (0,3,5) 4.38 8.75 ± 0.19 3.06 ± 0.06 8.22 ± 1.15 4.84 ± 0.73 5.71 ± 0.26 5.15 ± 0.10 (2,1,0) 0.53 2.59 ± 0.08 1.60 ± 0.05 2.37 ± 0.15 6.04 ± 0.41 0.76 ± 0.02 1.03 ± 0.54 (4,2,0) 0.92 2.16 ± 0.05 2.47 ± 0.07 8.54 ± 0.74 4.23 ± 0.46 0.79 ± 0.00 0.95 ± 0.09 (5,3,0) 1.17 2.17 ± 0.04 2.56 ± 0.08 9.04 ± 0.92 3.87 ± 0.10 0.86 ± 0.02 1.35 ± 0.56 SP ($\infty$,6,4) 32.57 9.95 ± 0.52 1.24 ± 0.02 11.16 ± 8.81 9.83 ± 5.92 3.13 ± 0.38 11.14 ± 12.06 ($\infty$,4,3) 19.84 3.29 ± 0.10 5.87 ± 0.16 0.79 ± 0.54 4.94 ± 0.62 3.58 ± 0.45 2.96 ± 4.58 (4,6$\infty$) 0.65 10.53 ± 0.48 9.95 ± 0.52 5.84 ± 0.34 5.90 ± 1.76 4.89 ± 0.13 1.51 ± 0.72 (3,4$\infty$) 0.54 3.38 ± 0.11 5.87 ± 0.16 1.01 ± 0.19 1.38 ± 0.46 0.83 ± 0.01 0.51 ± 0.02 PA (0,5k,10k) 3.9 3.69 ± 0.76 3.93 ± 0.98 8.00 ± 0.60 9.27 ± 1.78 4.92 ± 0.58 4.22 ± 0.63 (0,10k,20k) 3.14 3.42 ± 0.59 6.38 ± 3.48 6.90 ± 1.46 9.03 ± 1.6 6.29 ± 0.87 3.43 ± 0.19 (0,15k,25k) 2.72 5.62 ± 0.40 2.48 ± 0.03 8.00 ± 0.78 9.07 ± 2.43 8.98 ± 1.10 3.57 ± 0.74 (10k,5k,0) 7.4 3.09 ± 0.15 3.15 ± 0.16 10.56 ± 0.73 14.43 ± 4.45 1.52 ± 0.05 4.88 ± 0.90 (20k,10k,0) 5.81 2.50 ± 0.23 2.55 ± 0.16 7.74 ± 0.46 8.43 ± 3.46 1.38 ± 0.04 4.50 ± 1.10 (25k,15k,0) 5.08 2.79 ± 0.28 2.37 ± 0.02 7.82 ± 0.36 6.27 ± 3.87 1.39 ± 0.06 2.38 ± 0.73
Table 11. ogbl-ddi results reported in MRR with the best bolded and the second best underlined.
Split Models RA GCN BUDDY NCNC LPFormer NeoGNN SEAL CN (0,1,2) 4.8 5.34 ± 0.24 1.65 ± 0.17 0.79 ± 0.20 0.82 ± 0.21 0.92 ± 0.01 0.89 ± 0.39 (0,2,4) 4.25 4.88 ± 0.41 2.75 ± 0.16 1.39 ± 0.81 0.85 ± 0.14 2.12 ± 0.77 1.17 ± 0.32 (0,3,5) 3.86 3.48 ± 0.26 2.15 ± 0.08 1.88 ± 0.57 1.06 ± 0.17 1.75 ± 0.44 0.98 ± 0.37 (2,1,0) 0.46 1.14 ± 0.20 0.78 ± 0.06 2.05 ± 2.58 3.13 ± 0.63 0.94 ± 0.40 0.87 ± 0.49 (4,2,0) 0.59 1.01 ± 0.20 1.09 ± 0.22 0.61 ± 0.17 3.46 ± 0.61 2.66 ± 1.94 0.62 ± 0.08 (5,3,0) 0.54 1.41 ± 0.32 1.40 ± 0.35 0.58 ± 0.03 3.94 ± 0.28 2.69 ± 1.89 0.62 ± 0.07 PA (0,5k,10k) 5.29 2.61 ± 0.38 1.81 ± 0.35 3.44 ± 1.73 1.19 ± 0.11 1.96 ± 0.72 0.83 ± 0.30 (0,10k,20k) 3.6 1.96 ± 0.22 1.74 ± 0.28 2.70 ± 0.53 0.84 ± 0.21 1.31 ± 0.30 1.35 ± 0.87 (0,15k,25k) 2.54 1.91 ± 0.14 1.92 ± 0.08 2.33 ± 1.18 1.46 ± 0.31 1.31 ± 0.20 1.48 ± 0.69 (10k,5k,0) 7.06 4.68 ± 1.19 1.19 ± 0.32 12.55 ± 15.18 14.19 ± 5.59 3.63 ± 0.50 4.64 ± 4.21 (20k,10k,0) 1.5 2.04 ± 0.77 2.71 ± 1.60 10.03 ± 9.20 8.26 ± 1.13 2.89 ± 0.81 1.31 ± 0.92 (25k,15k,0) 1.68 1.14 ± 0.39 2.09 ± 0.72 4.66 ± 4.89 8.51 ± 0.82 1.58 ± 0.98 1.71 ± 0.56
Appendix G Generalization Results
This section includes Tables 12 to 14, displaying results for the graph-specific generalization methods conducted on the LPShift dataset splits.
Table 12. ogbl-collab (MRR) results under LPShift after applying graph-specific generalization methods.
CN (0,1,2) (0,2,4) (0,3,5) (2,1,0) (4,2,0) (5,3,0) DropEdge 15.54 ± 0.98 16.16 ± 0.17 16.34 ± 0.17 2.61 ± 0.09 2.88 ± 0.11 4.99 ± 0.08 TC 11.27 ± 2.03 12.39 ± 1.43 14.99 ± 1.69 5.28 ± 0.05 5.23 ± 0.05 6.03 ± 0.10 GCN+BUDDY 8.50 ± 1.10 12.85 ± 0.83 15.35 ± 0.96 5.46 ± 0.11 5.36 ± 0.12 5.96 ± 0.06 RA+BUDDY 3.94 ± 0.51 6.61 ± 0.30 7.16 ± 0.09 4.32 ± 0.19 4.45 ± 0.12 4.93 ± 0.11 SP ($\infty$,6,4) ($\infty$,4,3) (4,6$\infty$) (3,4$\infty$) DropEdge 12.31 ± 0.51 17.11 ± 1.02 5.34 ± 0.43 2.93 ± 0.18 TC 7.25 ± 0.81 9.33 ± 1.66 6.88 ± 0.30 6.24 ± 0.13 GCN+BUDDY 7.38 ± 0.82 9.60 ± 0.39 6.86 ± 1.13 6.47 ± 0.24 RA+BUDDY 3.84 ± 0.59 7.63 ± 0.25 6.86 ± 1.13 6.47 ± 0.24 PA (0,50,100) (0,100,200) (0,150,250) (100,50,0) (200,100,0) (250,150,0) DropEdge 21.35 ± 0.36 13.84 ± 0.64 12.85 ± 0.78 26.09 ± 0.62 15.68 ± 0.85 13.13 ± 0.94 TC 18.82 ± 1.35 12.13 ± 1.04 10.63 ± 0.55 15.84 ± 1.13 9.15 ± 0.39 6.70 ± 0.20 GCN+BUDDY 15.92 ± 1.01 9.47 ± 0.31 9.60 ± 0.41 14.34 ± 1.05 8.35 ± 0.34 5.50 ± 0.33 RA+BUDDY 15.92 ± 1.01 9.47 ± 0.31 9.60 ± 0.41 14.34 ± 1.05 8.35 ± 0.34 5.50 ± 0.33
Table 13. ogbl-ppa (MRR) results under LPShift after applying graph-specific generalization methods.
CN (0,1,2) (0,2,4) (0,3,5) (2,1,0) (4,2,0) (5,3,0) DropEdge 7.83 ± 0.27 3.83 ± 0.25 3.06 ± 0.06 1.61 ± 0.04 2.47 ± 0.07 2.56 ± 0.08 TC 5.27 ± 0.34 2.91 ± 0.06 2.67 ± 0.13 3.44 ± 0.08 3.45 ± 0.10 3.55 ± 0.13 GCN+BUDDY 4.48 ± 0.33 3.79 ± 0.28 3.16 ± 0.10 3.19 ± 0.08 3.25 ± 0.09 3.36 ± 0.13 RA+BUDDY 4.04 ± 0.26 3.42 ± 0.20 2.95 ± 0.13 2.58 ± 0.09 3.04 ± 0.05 3.10 ± 0.15 SP ($\infty$,6,4) ($\infty$,4,3) (4,6$\infty$) (3,4$\infty$) DropEdge 3.86 ± 0.39 5.87 ± 0.16 3.86 ± 0.39 5.87 ± 0.16 TC 4.00 ± 0.29 4.82 ± 0.39 13.2 ± 0.45 2.89 ± 0.09 GCN+BUDDY 4.00 ± 0.20 5.53 ± 0.92 14.41 ± 0.67 2.93 ± 0.15 RA+BUDDY 3.89 ± 0.23 5.53 ± 0.94 13.63 ± 0.97 2.51 ± 0.14 PA (0,5k,10k) (0,10k,20k) (0,15k,25k) (10k,5k,0) (20k,10k,0) (25k,15k,0) DropEdge 3.93 ± 0.98 6.38 ± 3.48 2.49 ± 0.01 3.13 ± 0.10 2.56 ± 0.19 2.40 ± 0.03 TC 3.62 ± 0.21 3.13 ± 0.12 3.33 ± 1.38 1.78 ± 0.16 1.50 ± 0.0008 1.53 ± 0.04 GCN+BUDDY 3.66 ± 0.38 3.19 ± 0.22 2.88 ± 0.06 2.05 ± 0.05 1.67 ± 0.03 1.66 ± 0.02 RA+BUDDY 3.66 ± 0.38 3.19 ± 0.22 2.88 ± 0.06 2.05 ± 0.04 1.65 ± 0.05 1.64 ± 0.03
Table 14. ogbl-ddi (MRR) results under LPShift after applying graph-specific generalization methods.
CN (0,1,2) (0,2,4) (0,3,5) (2,1,0) (4,2,0) (5,3,0) DropEdge 1.65 ± 0.17 2.75 ± 0.16 2.15 ± 0.08 0.78 ± 0.06 1.09 ± 0.22 1.40 ± 0.35 TC 1.10 ± 0.01 1.01 ± 0.01 0.94 ± 0.01 1.04 ± 0.02 0.90 ± 0.02 0.98 ± 0.03 GCN+BUDDY 0.87 ± 0.21 1.45 ± 0.35 1.65 ± 0.22 1.20 ± 0.14 2.44 ± 0.27 3.19 ± 0.36 RA+BUDDY 0.87 ± 0.21 1.45 ± 0.35 1.65 ± 0.22 1.20 ± 0.14 2.44 ± 0.27 3.19 ± 0.36 PA (0,5k,10k) (0,10k,20k) (0,15k,25k) (10k,5k,0) (20k,10k,0) (25k,15k,0) DropEdge 1.81 ± 0.35 1.74 ± 0.28 1.92 ± 0.08 3.83 ± 3.22 2.71 ± 1.60 11.07 ± 1.44 TC 1.04 ± 0.01 1.06 ± 0.02 1.24 ± 0.02 1.00 ± 0.03 1.66 ± 0.01 1.95 ± 0.02 GCN+BUDDY 1.26 ± 0.28 1.43 ± 0.10 1.04 ± 0.23 9.16 ± 9.27 9.84 ± 1.97 9.50 ± 2.18 RA+BUDDY 1.26 ± 0.28 1.43 ± 0.11 1.04 ± 0.23 9.16 ± 9.27 9.84 ± 1.97 9.50 ± 2.18
Appendix H Traditional OOD Generalization Method Results
This section includes table detailing the performance of domain-agnostic/traditional OOD generalization methods on all of LPShift’s tested dataset splits.
Table 15. MRR Results for traditional OOD generalization methods on LPShift’s ogbl-collab.
Models CN Splits SP Splits PA Splits (0, 1, 2) (0, 2, 4) (0, 3, 5) ($\infty$, 6, 4) ($\infty$, 4, 3) (0, 50, 100) (0, 100, 200) (0, 150, 250) GCN 14.89 ± 0.67 17.75 ± 0.44 19.08 ± 0.26 12.64 ± 0.85 15.50 ± 0.16 19.81 ± 0.59 12.68 ± 0.20 12.73 ± 0.32 +IRM 7.40 ± 0.32 6.97 ± 0.55 5.50 ± 0.71 5.84 ± 0.58 5.28 ± 0.17 11.18 ± 1.46 6.88 ± 1.65 6.95 ± 1.09 +VREx 14.31 ± 0.76 16.95 ± 0.22 18.86 ± 0.13 13.17 ± 0.58 15.75 ± 0.50 21.69 ± 0.18 15.23 ± 0.44 14.66 ± 0.14 +GroupDRO 5.13 ± 0.55 5.19 ± 0.61 5.51 ± 0.67 6.27 ± 0.91 6.36 ± 0.53 9.56 ± 1.93 7.39 ± 0.65 6.04 ± 0.90 +DANN 15.69 ± 0.28 16.83 ± 0.18 18.80 ± 0.20 14.01 ± 0.37 15.60 ± 0.45 21.77 ± 0.31 15.61 ± 0.30 14.89 ± 0.11 +Deep CORAL 14.31 ± 0.76 16.95 ± 0.22 18.84 ± 0.10 13.26 ± 0.68 15.60 ± 0.29 21.69 ± 0.18 15.00 ± 0.28 14.59 ± 0.13 (2, 1, 0) (4, 2, 0) (5, 3, 0) (4, 6, $\infty$) (3, 4, $\infty$) (100, 50, 0) (200, 100, 0) (250, 150, 0) GCN 5.04 ± 0.13 4.87 ± 0.17 6.78 ± 0.19 4.75 ± 0.10 3.81 ± 0.16 24.33 ± 0.53 13.97 ± 0.52 12.98 ± 0.45 +IRM 2.42 ± 0.28 3.21 ± 0.29 3.08 ± 0.41 2.39 ± 0.37 1.81 ± 1.11 2.72 ± 1.12 2.00 ± 0.90 1.99 ± 0.70 +VREx 5.10 ± 0.04 5.60 ± 0.25 6.91 ± 0.15 4.67 ± 0.08 3.92 ± 0.22 23.06 ± 0.59 12.53 ± 0.21 11.81 ± 0.31 +GroupDRO 3.38 ± 0.26 2.77 ± 0.07 2.99 ± 0.02 3.04 ± 0.75 2.88 ± 0.74 3.49 ± 0.64 3.32 ± 0.42 3.49 ± 0.92 +DANN 5.47 ± 0.11 5.40 ± 0.25 6.98 ± 0.28 8.63 ± 0.32 6.65 ± 0.52 21.83 ± 0.61 12.53 ± 0.46 11.67 ± 0.22 +Deep CORAL 5.13 ± 0.08 5.49 ± 0.13 6.82 ± 0.13 8.47 ± 0.18 6.72 ± 0.36 23.10 ± 0.67 12.52 ± 0.20 11.80 ± 0.31
Table 16. MRR Results for traditional OOD generalization methods on LPShift’s ogbl-ppa.
Models CN Splits SP Splits PA Splits (0, 1, 2) (0, 2, 4) (0, 3, 5) ($\infty$, 6, 4) ($\infty$, 4, 3) (0, 5k, 10k) (0, 10k, 20k) (0, 15k, 25k) GCN 7.81 ± 0.08 8.18 ± 0.12 8.75 ± 0.19 6.34 ± 0.72 5.05 ± 0.08 3.69 ± 0.76 3.42 ± 0.59 5.62 ± 0.40 +IRM 1.99 ± 0.45 1.97 ± 1.37 2.30 ± 1.05 6.44 ± 0.62 4.98 ± 0.15 3.01 ± 0.61 5.24 ± 0.20 3.13 ± 0.56 +VREx 6.91 ± 0.18 7.32 ± 0.25 7.81 ± 0.21 6.35 ± 0.83 5.19 ± 0.23 4.30 ± 0.77 3.94 ± 0.83 4.71 ± 0.85 +GroupDRO 2.34 ± 0.19 2.31 ± 0.27 2.52 ± 0.37 5.03 ± 1.91 2.62 ± 1.03 3.06 ± 0.44 4.70 ± 0.60 2.77 ± 0.32 +DANN 7.06 ± 0.19 7.34 ± 0.14 7.82 ± 0.18 6.34 ± 1.05 5.30 ± 0.38 4.28 ± 0.87 4.74 ± 1.12 5.66 ± 0.36 +Deep CORAL 6.91 ± 0.18 7.32 ± 0.25 7.79 ± 0.22 6.25 ± 0.62 5.18 ± 0.23 4.19 ± 0.76 4.08 ± 0.73 4.57 ± 0.74 (2, 1, 0) (4, 2, 0) (5, 3, 0) (4, 6, $\infty$) (3, 4, $\infty$) (10k, 5k, 0) (20k, 10k, 0) (25k, 15k, 0) GCN 2.59 ± 0.08 2.16 ± 0.05 2.17 ± 0.04 9.95 ± 0.52 3.29 ± 0.10 3.09 ± 0.15 2.50 ± 0.23 2.79 ± 0.28 +IRM 1.78 ± 0.56 3.05 ± 0.17 3.00 ± 0.32 2.34 ± 2.07 1.88 ± 0.64 3.98 ± 0.23 3.21 ± 0.28 3.09 ± 0.12 +VREx 3.51 ± 0.28 2.91 ± 0.11 2.87 ± 0.17 9.42 ± 0.85 3.48 ± 0.32 3.41 ± 0.86 2.50 ± 0.48 2.57 ± 0.49 +GroupDRO 1.85 ± 0.42 3.20 ± 0.23 3.23 ± 0.16 2.19 ± 0.29 2.09 ± 0.60 4.42 ± 0.26 3.37 ± 0.08 3.07 ± 0.13 +DANN 3.62 ± 0.29 2.72 ± 0.18 2.73 ± 0.11 9.69 ± 0.46 3.32 ± 0.07 2.71 ± 0.25 2.13 ± 0.18 2.20 ± 0.20 +Deep CORAL 3.43 ± 0.34 2.93 ± 0.11 2.70 ± 0.08 9.42 ± 0.85 3.48 ± 0.32 2.57 ± 0.45 2.13 ± 0.36 2.14 ± 0.15
Table 17. MRR Results for traditional OOD generalization methods on LPShift’s ogbl-ddi.
Models CN Splits PA Splits (0, 1, 2) (0, 2, 4) (0, 3, 5) (0, 5k, 10k) (0, 10k, 20k) (0, 15k, 25k) GCN 5.34 ± 0.24 4.88 ± 0.41 3.48 ± 0.26 2.61 ± 0.38 1.96 ± 0.22 1.91 ± 0.14 +IRM 0.77 ± 0.16 1.36 ± 0.46 1.26 ± 0.19 1.34 ± 0.43 1.55 ± 0.57 1.90 ± 1.17 +VREx 0.64 ± 0.02 0.61 ± 0.02 1.41 ± 0.43 2.09 ± 0.29 1.28 ± 0.71 1.13 ± 0.21 +GroupDRO 0.64 ± 0.02 1.89 ± 0.72 0.77 ± 0.16 1.12 ± 0.14 0.84 ± 0.22 2.22 ± 1.11 +DANN 0.63 ± 0.02 0.98 ± 0.82 1.44 ± 0.50 2.19 ± 0.16 1.12 ± 0.54 1.19 ± 0.20 +Deep CORAL 0.64 ± 0.02 0.61 ± 0.02 1.29 ± 0.30 2.22 ± 0.44 1.23 ± 0.72 1.16 ± 0.20 (2, 1, 0) (4, 2, 0) (5, 3, 0) (10k, 5k, 0) (20k, 10k, 0) (25k, 15k, 0) GCN 1.14 ± 0.20 1.01 ± 0.20 1.41 ± 0.32 4.68 ± 1.19 2.04 ± 0.77 1.14 ± 0.39 +IRM 4.32 ± 0.64 4.03 ± 1.02 4.94 ± 2.16 5.22 ± 2.91 5.69 ± 2.68 6.74 ± 0.85 +VREx 4.08 ± 0.63 3.63 ± 0.01 4.14 ± 0.97 3.67 ± 1.76 2.61 ± 2.36 6.27 ± 0.58 +GroupDRO 4.58 ± 1.33 3.68 ± 0.31 3.62 ± 0.01 4.13 ± 3.55 3.05 ± 3.23 6.37 ± 0.03 +DANN 4.91 ± 1.31 3.66 ± 0.10 4.78 ± 1.48 4.11 ± 2.44 2.22 ± 2.03 6.29 ± 1.32 +Deep CORAL 4.08 ± 0.63 3.63 ± 0.01 4.06 ± 1.00 3.51 ± 1.63 1.71 ± 1.71 6.19 ± 0.59
Appendix I Earth Mover’s Distance (EMD) Results
This section contains the EMD calculations for all dataset splits and a visualization for the EMD calculations on the ”Forward” ogbl-ddi dataset split.
Figure 8. The EMD values calculated between the heuristic scores of training and testing samples on the ”Backward” LPShift splits before and after applying structural generalization methods. Note: The tested heuristics correspond to their labelled LPShift splits, so as to simulate the dataset splitting.
[…]
The EMD values calculated between the heuristic scores of training and testing samples on the ”Backward” LPShift splits before and after applying structural generalization methods. Note: The tested heuristics correspond to their labelled LPShift splits, so as to simulate the dataset splitting.
Table 18. EMD calculations for ogbl-collab, ogbl-ppa, and ogbl-ddi on the forward and backward splits. (Note: Scores with a distance multiple-times different than the baseline are in bold for ogbl-collab.)
Split ogbl-collab ogbl-ppa ogbl-ddi Baseline DropEdge EPS Baseline DropEdge EPS Baseline DropEdge EPS CN (0, 1, 2) 1.31 1.31 3.6 2.82 2.82 $>$24hrs 3.87 3.68 2.47 (0, 2, 4) 1.6 1.6 2.52 3.13 3.13 $>$24hrs 4.29 4.1 2.09 (0, 3, 5) 1.45 1.65 2.22 3.05 3.19 $>$24hrs 4.44 4.27 2.05 (2, 1, 0) 1.87 1.15 2.91 3.1 2.36 $>$24hrs 5.81 5.62 4.1 (4, 2, 0) 2.26 1.49 2.99 3.3 2.55 $>$24hrs 5.81 5.62 4.17 (5, 3, 0) 2.28 1.52 2.14 3.19 2.44 $>$24hrs 5.69 5.44 4.08 SP ($\infty$, 6, 4) 5.93 5.94 0.012 5.81 5.84 $>$24hrs ($\infty$, 4, 3) 5.35 5.38 0.003 1.36 1.4 $>$24hrs (4, 6, $\infty$) 3.6 3.53 1.22 2.14 2.14 $>$24hrs (3, 4, $\infty$) 1.85 1.78 3.23 0.72 0.72 $>$24hrs PA (0, 50, 100) 1.87 1.89 3.42 2.55 2.55 $>$24hrs 4.33 4.35 0.49 (0, 100, 200) 2.29 2.32 2.72 2.76 2.76 $>$24hrs 4.24 4.25 0.19 (0, 150, 250) 2.34 2.36 3.08 2.78 2.78 $>$24hrs 4.14 4.16 0.20 (100, 50, 0) 4.29 4.3 2.48 2.96 2.96 $>$24hrs 6.85 6.92 0.71 (200, 100, 0) 3.79 3.82 0.78 2.68 2.68 $>$24hrs 6.41 6.49 0.58 (250, 150, 0) 3.48 3.5 2.48 2.48 2.48 $>$24hrs 5.98 6.02 0.44
Appendix J Dataset Licenses
The dataset splitting strategy proposed in this paper is built using Pytorch Geometric (PyG). As such, this project’s software and the PyG datasets are freely-available under the MIT license.
Appendix K Limitations
The proposed dataset splitting strategy is restricted to inducing distribution shifts solely with neighborhood heuristics on static graphs. So, it does not directly consider other types of possible distribution shifts for the link prediction task (i.e. spatio-temporal (Zhang et al., 2022) or size (Zhou et al., 2022a) shift). Additionally, since the neighborhood heuristics compute discrete scores produced from an input graph’s structural information and effectively training GNN4LP models requires no leakage with validation/testing, it may be difficult to determine the correct thresholds to extract a meaningful number of samples. For Common Neighbors and Preferential-Attachment, this is especially relevant with smaller training graphs, given that larger and/or denser graphs have inherently more edges. Therefore, larger and denser graphs have inherently more possible Common Neighbors and Preferential-Attachment scores. For Shortest-Path, splitting can be exceptionally difficult for denser graphs, as demonstrated with the tiny split sizes for ogbl-ppa in Table 7.
Appendix L Impact Statement
Our proposed dataset splitting strategy mimics the formatting of PyTorch Geometric datasets. This means that our strategy is simple to implement, enabling future work involved with understanding this type of structural shift for link prediction and promoting beginner-friendly practices for artificial intelligence research. Additionally, since the structural shift we propose in this article affects real-life systems, which integrate link prediction models, this research can provide a foundation for the improvement of relevant technologies; which holds positive ramifications for society and future research. No apparent risk is related to the contribution of this work.